A couple with a new baby girl want to set aside money today that will cover her four years of college. This will be a ONE-time contribution to a mutual fund that returns 9.00% per year on average. Currently, college tuition costs $39,200.00 for four years, but it is growing at 2.00% per year. The "baby" girl will start college in 18.00 years.

How much money must they set aside today to cover the cost of college?
Answer #1
Step 1 - Find out the Future value of college tuition cost at the end of 18th year
We can use the Future value of sum formula to calculate this amount.
Future value of sum = P * (1+r)^n
Future value of sum = College tuition cost at the end of 18th year = ?
P = Current college tuition cost = $39200
r = growth rate in cost = 2%
n = number of years = 18
Future value of sum = 39200 * (1+0.02)^18
Future value of sum = 55987.25
College Tuition cost at the end of 18th Year = $55,987.25
Step 2 - Find out the present value of college tuition cost i.e.$55,987.25
We can use the present value of sum formula to know the amount need to set out today to cover the cost of college
Present value of sum = F/(1+r)^n
Present value of sum = amount invested today in mutual fund = ?
F = College Tuition cost at the end of 18th Year = $55,987.25
r = mutual fund rate of return per year = 9%
n = number of years = 18
Present value of sum = 55987.25/(1+0.09)^18
Present value of sum = 11868.95
Couple must set aside $11,868.95 today to cover the cost of college.